### v2

# Fine-tuning Diffusion Policies with Backpropagation Through Diffusion Timesteps

Abstract
Diffusion policies, widely adopted in decision-making scenarios such as robotics, gaming and autonomous driving, are capable of learning diverse skills from demonstration data due to their high representation power. However, the sub-optimal and limited coverage of demonstration data could lead to diffusion policies that generate sub-optimal trajectories and even catastrophic failures. While reinforcement learning (RL)-based fine-tuning has emerged as a promising solution to address these limitations, existing approaches struggle to effectively adapt Proximal Policy Optimization (PPO) to diffusion models. This challenge stems from the computational intractability of action likelihood estimation during the denoising process, which leads to complicated optimization objectives. In our experiments starting from randomly initialized policies, we find that online tuning of Diffusion Policies demonstrates much lower sample efficiency compared to directly applying PPO on MLP policies (MLP+PPO). To address these challenges, we introduce NCDPO, a novel framework that reformulates Diffusion Policy as a noise-conditioned deterministic policy. By treating each denoising step as a differentiable transformation conditioned on pre-sampled noise, NCDPO enables tractable likelihood evaluation and gradient backpropagation through all diffusion timesteps. Our experiments demonstrate that NCDPO achieves sample efficiency comparable to MLP+PPO when training from scratch, outperforming existing methods in both sample efficiency and final performance across diverse benchmarks, including continuous robot control and multi-agent game scenarios. Furthermore, our experimental results show that our method is robust to the number denoising timesteps.
1 Introduction
Recently, diffusion models have been widely adopted as policy classes in decision-making scenarios such as robotics  [5, 21, 2, 33, 27, 15], gaming [15, 34], and autonomous driving  [13, 30]. Although Diffusion Policies have shown remarkable capabilities in learning diverse behaviors from demonstration data [5], Diffusion Policy could show sub-optimal performance when the demonstration data is sub-optimal or only covers a limited set of environment states. To further optimize the performance of pretrained policies, Reinforcement Learning (RL) is adopted as a natural choice for fine-tuning pre-trained Diffusion Policies through interaction with the environment.
Currently, the most effective approach, DPPO (Diffusion Policy Policy Optimization[20] employs Policy Gradient (PG) approaches to enhance the performance of pre-trained Diffusion Policy in continuous control tasks. By treating the denoising process of Diffusion Policy as a low-level Markov Decision Process, DPPO optimizes the Gaussian likelihood of all denoising steps. However, through our extensive experiments, we find fine-tuning Diffusion Policies with RL faces a challenge of sample efficiency. Specifically, in our RL experiments starting from randomly initialized policies, we find that training Diffusion Policy with DPPO could lead to worse sample efficiency and final performance than training an MLP policy with standard RL. We hypothesize that the training efficiency gap occurs since DPPO uses a much larger action space for RL training, which impedes the sample efficiency of RL training. Therefore, a question becomes particularly important: Can we design a more effective fine-tuning approach for Diffusion Policy that avoids enlarging the action space during RL training?
In this work, we present Noise-Conditioned Diffusion Policy Optimization (NCDPO), a sample-efficient RL algorithm for fine-tuning Diffusion Policies. NCDPO formulates the denoising process of Diffusion Policy as a noise-conditioned inference process, ensuring the RL objective only contain the likelihood of the interactive actions, i.e. actions generated by Diffusion Policy to interact with the environment. In the policy update phase, the gradients with respect to the policy parameters are computed with Backpropagation through Diffusion Timesteps (BPDT). When performing RL training on randomly initialized policies, we show that training Diffusion Policy with NCDPO achieves comparable sample efficiency with training an MLP policy with RL.
In summary, our main contribution is NCDPO, a novel framework which is applicable to both continuous and discrete environments, to fine-tune Diffusion Policies, by formulating denoising steps as deterministic generation process and apply PPO. We also evaluate NCDPO on a set of environments, ranging from continuous robot control and multi-agent coordination tasks. We demonstrate that NCDPO obtains higher sample efficiency and stronger final performance than baseline methods across all evaluated environments. Finally, our ablation study reveals that that NCDPO is robust to the number of diffusion timesteps and remains highly sample efficient when the number of diffusion timesteps is large.
2 Related Work
Diffusion Models and Diffusion Policies.
Diffusion-based generative models have demonstrated remarkable effectiveness in the domains of visual content generation [23, 26, 19].One central capability of Diffusion Models is the denoising process that iteratively refines sampled noises into clean datapoints. [9, 24, 25]. Beyond their success in content generation, diffusion models have increasingly been adapted for decision-making tasks across a range of domains, including robotics [5, 21, 2, 33, 27, 15], autonomous driving [13, 30], and gaming [15, 34]. In robotics, most existing work trains Diffusion Policies through imitation learning. For instance, Reuss et al. [21] predict future action chunks using goal-conditioned imitation learning, while [33] integrate Diffusion Policies with compact 3D representations extracted from point clouds. To further enhance the quality of generated behaviors, return signal or goal conditioning is applied to encourage the generation of high-value actions [10, 1, 12].
Fine-tuning Diffusion Policy with Reinforcement Learning.
Recent works have aimed to enhance learned Diffusion Policy through fine-tuning with Reinforcement Learning approaches. A line of work has been focusing on integrating Diffusion Policies with Q-learning using offline data [4, 11, 28, 7, 22, 35, 18]. In addition to offline reinforcement learning, recent advancements have explored fine-tuning Diffusion Policies with online RL algorithms, for example, aligning the score function with the action gradient [31], or employing the diffusion model as a policy extraction mechanism within implicit Q-learning [8]. Most recently, [20] formulates the denoising process of Diffusion Policy as a "Diffusion MDP", enabling the application of RL algorithms to optimize all denoising steps with online feedback. In this work, we investigate an alternative representation for the denoising process that enables sample efficient fine-tuning of Diffusion Policy.
3 Preliminary
Markov Decision Process. A Markov Decision Process (MDP) is defined as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P_{0},P,R,\gamma\rangle$ where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ is the action space, $P_{0}$ is the distribution of initial states, $P$ is the transition function, $R$ is the reward function and $\gamma$ is the discount factor. At timestep $t$, a policy $\pi$ generates an action $a_{t}\in\mathcal{A}$ at state $s_{t}$. The goal is to find a policy $\pi$ that maximizes the objective of expected discounted return,
$J(\pi)=\mathbb{E}_{s_{t},a_{t}}[\sum_{t\geq 0}\gamma^{t}R(s_{t},a_{t})]$ (1)
Proximal Policy Optimization (PPO). PPO is a reinforcement learning approach that optimizes the policy by estimating the policy gradient. In each iteration, given the last iteration policy $\pi_{\theta_{k}}$, PPO maximizes the clipped objective,
$\begin{split}L(\theta|\theta_{k})=\mathbb{E}_{\tau}&\Biggl{[}\sum_{t}\min\Biggl{(}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{k}}(a_{t}|s_{t})}A^{\pi_{\theta_{k}}}(s_{t},a_{t}),\\[2.84526pt] &\text{clip}\Bigl{(}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{k}}(a_{t}|s_{t})},1-\epsilon,1+\epsilon\Bigr{)}A^{\pi_{\theta_{k}}}(s_{t},a_{t})\Biggr{)}\Biggr{]}\end{split}$ (2)
where $A^{\pi_{\theta_{k}}}(s_{t},a_{t})$ is the estimated advantage for action $a_{t}$ at state $s_{t}$.
Diffusion Policy. Diffusion Policy $\pi_{\theta}$ is a diffusion model that generates actions $a$ by conditioning on states $s$. In Diffusion Policy training, the forward process gradually adds Gaussian noise to the training data to obtain a chain of noisy datapoints $a^{0},a^{1},\dots,a^{K}$,
$q(a^{1:K}|a^{0}):=\prod_{k=1}^{K}q(a^{k}|a^{k-1}),\quad\quad q(a^{k}|a^{k-1}):=\mathcal{N}(a^{k};\sqrt{1-\beta_{k}}a^{k-1},\beta_{k}I)$ (3)
Diffusion Policy could generate actions with a reverse process or denoising process that gradually denoises a Gaussian noise $a^{K}\sim\mathcal{N}(a^{K};0,I)$ with learned Gaussian transitions,
$\pi_{\theta}(a^{0:K}|s):=\prod_{k=1}^{K}\pi_{\theta}(a^{k-1}|a^{k},s),\quad\quad\pi_{\theta}(a^{k-1}|a^{k},s):=\mathcal{N}(a^{k-1}|\mu_{\theta}(a^{k},k,s),\sigma^{2}_{k}I)$ (4)
where $\sigma$ is a fixed noise schedule for action generation, $\beta$ denotes the forward process variances and is held as constant, and $\theta$ is the parameter of Diffusion Policy. To avoid ambiguity, we use interactive actions to denote the action $a^{0}$ that is used for interacting with the environment and latent actions to denote actions $a^{1},\cdots,a^{K}$ that are generated during the denoising process. For more training details on diffusion models, please refer to [9].
Diffusion Policy Policy Optimization (DPPO). Note that the action likelihood $\pi_{\theta}(a_{t}^{0}|s_{t})$ of Diffusion Policy $\pi_{\theta}$ is intractable,
[…]
The intractability of the action likelihood makes it impossible to directly fine-tune Diffusion Policy with PPO since the RL loss (Eq. 2) requires computing the exact action likelihood. To address this challenge, DPPO [20] proposes to formulate the denoising process as a low-level "Diffusion MDP" $\mathcal{M}_{\text{Diff}}$. In $\mathcal{M}_{\text{Diff}}$, a state is defined as a combination of the environment state and a latent action $\hat{s}^{k}_{t}=(s_{t},a^{k}_{t})$. For $k=K,\cdots,1$, the transition from $\hat{s}^{k}_{t}$ to $\hat{s}^{k-1}_{t}$ represents the denoising process and takes no actual change on the environment state. For a denoising step $k\in[1,K]$, the state $\hat{s}^{k}_{t}=(s_{t},a^{k}_{t})$ transits to $\hat{s}^{k-1}_{t}=(s_{t},a^{k-1}_{t})$. After the denoising process is finished at $k=0$, the interactive action $a^{0}_{t}$ is used to interact with the environment and triggers the environment transition, i.e. the next state would be $\hat{s}_{t+1}^{K}=(s_{t+1},a_{t+1}^{K})$ where $s_{t+1}\sim P(s_{t},a^{0}_{t})$ and $a_{t+1}^{K}\sim N(0,I)$ is a newly generated Gaussian noise.
4 Sample Efficiency Challenge for Diffusion Policy Fine-Tuning
In this section, we aim to investigate the sample efficiency of fine-tuning Diffusion Policy with RL. Specifically, we compare the sample efficiency of training Diffusion Policy using DPPO and training an MLP policy using standard PPO. Since our study only focuses on the RL process, both the MLP policy and Diffusion Policy are randomly initialized before RL training without performing any additional behavior cloning. For conciseness, we denote training Diffusion Policy with DPPO as DP+DPPO and training an MLP policy with PPO as MLP+PPO.
Our investigations are carried out on two OpenAI Gym locomotion tasks, Walker2D and Halfcheetah. The training curves of MLP+PPO and DP+DPPO are shown in Fig. 1. Although Diffusion Policy has more powerful representation power than MLP policy [5], our results here surprisingly show that DP+DPPO is less sample efficient than MLP+PPO and could only achieves sub-optimal performance.
Why does this efficiency gap occur? We hypothesize that the underlying reason is that, by employing a two-level MDP formulation, DPPO actually significantly lengthens the MDP horizons in RL training to contain both the interactive actions and latent actions. This lengthened MDP horizon then results in difficulty in proper credit assignment. This insight raises a critical question: Can we design an alternative RL algorithm for Diffusion Policy fine-tuning that avoids enlarging the action space?
Figure 1: RL training from randomly initialized policy on Walker2D and HalfCheetah. Results are averaged over three seeds. Training curves indicate that DP+DPPO is less sample efficient than MLP+PPO and only achieves sub-optimal performance. Our approach, NCDPO, could fine-tune Diffusion Policy with high sample efficiency.
5 Noise-Conditioned Diffusion Policy Optimization
As discussed in Sec. 4, fine-tuning Diffusion Policy with a two-level MDP formulation could lead to sub-optimal sample efficiency and final performance. In this section, we present a novel sample-efficient RL training method for Diffusion Policy, Noise-Conditioned Diffusion Policy Optimization (NCDPO). In Sec. 5.1, we show that NCDPO formulates the denoising process of Diffusion Policy as a noise-conditioned inference process. In Sec. 5.2, we show that NCDPO ensures PPO training operates on the same action space as the environment, without relying on optimizing action likelihood of latent actions.
5.1 Denoising Process as a Noise-Conditioned Inference Process
Noise-conditioned Action Generation.
We decouple the stochastic and deterministic components of the denoising process. The stochastic component encompasses all the random noises sampled during the denoising process. The deterministic component further operates on these sampled noises with the model $\mu_{\theta}$.
Formally, Eq. 4 can be equivalently represented as,
$a^{k-1}=\mu_{\theta}(a^{k},k,s)+\sigma_{k}\cdot z^{k}\;\;\text{ where\ }z^{k}\sim\mathcal{N}(0,I)$ (5)
A straitforward indication of Eq. 5 is that, in each denoising step, the only stochastic component is the Gaussian noise $z^{k}$, while the computation of $\mu_{\theta}(a^{k},k,s)$ and addition between $\mu_{\theta}(a^{k},k,s)$ and $\sigma_{k}\cdot z^{k}$ are both deterministic. Therefore, the whole denoising process can be split into a noise sampling phase and a deterministic inference phase.
In the noise sampling phase, we generate a sequence of standard Gaussian noises $z^{1},\cdots,z^{K}$,
$z^{k}\sim\mathcal{N}(0,I)\;\;\text{ for }k=K,K-1,\dots,1$ (6)
$a^{K}\sim\mathcal{N}(0,I)$ (7)
In the deterministic inference phase, given noises $z^{1},\cdots,z^{K}$, $\mu_{\theta}$ is be used to compute the latent actions $a^{k}$ one by one. For $k=K,K-1,\dots,1$, $a^{k-1}$ is a linear combination of $\mu_{\theta}(a^{k},k,s)$ and $z^{k}$,
$a^{k-1}=\mu_{\theta}(a^{k},k,s)+\sigma_{k}\cdot z^{k}$ (8)
Consequently, the generated action $a^{0}$ can be computed by recursively applying Eq. 8,
$a^{0}$ $=\mu_{\theta}(\mu_{\theta}(\dots\mu_{\theta}(a^{K},K,s)\dots,2,s)+\sigma_{2}\cdot z^{2},1,s)+\sigma_{1}\cdot z^{1}$
$=f_{\theta}(s,a^{K},z^{1:K})$ (9)
MDP with Noise-augmented States.
As derived in Eq. 8 and Eq. 9, the denoising process can be partitioned into a noise sampling phase and a policy inference phase. We can incorporate the sampled noises into the MDP as part of the environment state. Formally, for the original environment MDP $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P_{0},P,R,\gamma\rangle$ we introduce MDP with noise-augmented states $\mathcal{M}_{noise}=\langle\mathcal{S}_{noise},\mathcal{A},P_{0},P_{noise},R,\gamma\rangle$. In $\mathcal{M}_{noise}$, each state $s_{noise}$ consists of a environment state $s\in\mathcal{S}$ and Gaussian noises $a^{K},z^{1},\cdots,z^{K}$. $\mathcal{M}_{noise}$ shares the same action space and reward function as the original MDP $\mathcal{M}$. In each decision-making step, the environment state transits to a new one, and the noises are all re-sampled.
(a) DPPO
(a) *
(b) NCDPO (ours)
(b) *
Figure 2: DPPO adopts a two-layer MDP design by combining the environment state with latent actions to form augmented states. In contrast, in each step, NCDPO first samples a group of random noises and computes the action based on the noises, resulting in a deterministic generation process (Eq. 9). Blue arrows in the figure indicate MDP transitions.
[.45]Refer to caption [.5]Refer to caption
Denoising Process as a Noise-Conditioned Policy.
Given a noise-augmented state $s_{noise}=(s,a^{K},z^{1:K})$, the deterministic inference phase of the denoising process can be represented as a Noise-Conditioned Policy $\pi^{NC}_{\theta}$ that generates the action $a^{0}$ using Eq. 9
Note that this noise-conditoned policy is a deterministic policy and could not be directly trained with PPO since the policy loss (Eq. 2) relies on a stochastic policy. Therefore, we introduce an additional operation to transform this deterministic policy into a stochastic one. Specifically, for continuous action space, we sample the final action $a^{0}$ near $f(a^{K},z^{1:K})$ from a Gaussian distribution with a learnable standard variation $\sigma_{act}$,
$\pi^{NC}_{\theta}(\cdot|z^{1:K},a^{K},s)=\mathcal{N}(f_{\theta}(s,a^{K},z^{1:K}),\sigma_{act}^{2})$ (10)
For discrete action space, we use softmax to sample the final action by treating $f_{\theta}(s,a^{K},z^{1:K})$ as logits,
$\pi^{NC}_{\theta}(a^{0}=i|z^{1:K},a^{K},s)\propto{\exp{(f_{\theta}(s,a^{K},z^{1:K})_{i}/T)}}$ (11)
where $i$ is the action index. $T$ is the temperature that allows the policy network to produce sharper action distributions.
5.2 Finetuning Noised-Conditioned Policy with PPO
Under the formulation of NCDPO, at each denoising timestep $t$, Gaussian noise $z^{t}$ is first sampled, and the action is then generated via Eq. 10 or Eq. 11. This way, we can apply PPO objective in Eq.12 , which utilizes a clipped objective to regularize updated policy from original policy, to optimize interactive action probabilities:
$\begin{split}L(\theta|\theta_{k})=\mathbb{E}_{\begin{subarray}{c}a_{t}^{0}\sim\pi^{NC}_{\theta_{k}}(z_{t}^{1:K},a_{t}^{K},s_{t})\\ \end{subarray}}&\Biggl{[}\sum_{t}\min\Biggl{(}\frac{\pi^{NC}_{\theta}(a_{t}^{0}|z_{t}^{1:K},a_{t}^{K},s_{t})}{\pi^{NC}_{\theta_{k}}(a_{t}^{0}|z_{t}^{1:K},a_{t}^{K},s_{t})}A^{\pi^{NC}_{\theta_{k}}}(a_{t}^{0}|s_{t}),\\[2.84526pt] &\text{clip}\Bigl{(}\frac{\pi^{NC}_{\theta}(a_{t}^{0}|z_{t}^{1:K},a_{t}^{K},s_{t})}{\pi^{NC}_{\theta_{k}}(a_{t}^{0}|z_{t}^{1:K},a_{t}^{K},s_{t})},1-\epsilon,1+\epsilon\Bigr{)}A^{\pi^{NC}_{\theta_{k}}}(a_{t}^{0}|s_{t})\Biggr{)}\Biggr{]}\end{split}$ (12)
As illustrated in Figure 16, in policy rollout process, each step begins by sampling a sequence of noises, which are then used by the Diffusion Policy to generate the corresponding action. These sampled noises are stored in the buffer. During training phase, the stored noises are reused to recompute the actions, enabling gradient backpropagation through the entire denoising process. This allows PPO to directly update all denoising steps of the diffusion policy.
Algorithm 1 NCDPO
0:  Noise-conditinoed policy $\pi^{NC}_{\theta}$, noise scheduler $\sigma$
1:  Parameters: $\gamma\in[0,1)$, $\varepsilon\in(0,1)$, $N_{\text{episodes}}$, $N_{\text{PPO}}$
2:  for $e=1,2,\ldots,N_{\text{episodes}}$ do
3:     $\text{buffer}\leftarrow\emptyset$
4:     for $t=0,1,2,\ldots,T-1$ do
5:        $a_{t}^{K},z_{t}^{1:K}\sim\mathcal{N}(0,I)$
6:        Sample $a_{t}^{0}$ from $\pi^{NC}(\cdot|z_{t}^{1:K},a_{t}^{K},s_{t})$
7:        $\log\pi^{NC}_{t}\leftarrow\pi^{NC}(a_{t}^{0}|z_{t}^{1:K},a_{t}^{K},s_{t})$
8:        Execute $a_{t}$, observe $r_{t}$, $s_{t+1}$
9:        $\text{buffer}\leftarrow\text{buffer}\cup\{s_{t},a_{t}^{K},r_{t},\log\pi_{t},z_{t}^{1:K}\}$
10:     end for
11:     for $\text{epoch}=1,2,\ldots,N_{\text{PPO}}$ do
12:        for mini-batch $b=1,2,\dots$ do
13:           Calculate PPO loss $L(\theta|\theta_{k})$ in Eq. 12, backpropagate gradients through diffusion timesteps and update parameter $\theta$
14:        end for
15:     end for
16:  end for
As Fig.2(a) shows, NCDPO models the denoising process as deterministic generation conditioned on pre-sampled noise $z^{1:K}_{t}$. During inference, interactive actions are obtained through recursive model inference in Eq. 9 and applying the action sampling step in Eq. 10 and Eq. 11.
6 Experiments
In this section, we provide a comprehensive evaluation of NCDPO across a variety of challenging environments. We begin by detailing the experimental setup in Sec. 6.1, followed by results on continuous robot control tasks in Sec. 6.2 and discrete multi-agent coordination tasks in Sec.6.3. Finally, we conduct ablation studies in Sec. 6.4 to assess the robustness and of NCDPO.
6.1 Environmental Setup
Environments: OpenAI Gym locomotion.
Our first set of experiments involves testing NCDPO on a series of well-established locomotion benchmarks from OpenAI Gym [3], namely: Hopper-v2, Walker2D-v2, and HalfCheetah-v2. The pre-trained Diffusion Policies used in these experiments are trained from the D4RL "medium" dataset [6], which contains a diverse range of pre-recorded trajectories. For the fine-tuning process, we use dense reward.
Environments: Robomimic.
We further evaluate the performance of NCDPO on robotic manipulation tasks within Robomimic benchmarks [14]. The specific scenarios we consider include Lift, Can, Square, and Transport, varying in difficulty. To ensure temporal consistency in actions, we employ action chunking with size 4 for Lift, Can, and Square, and size 8 for Transport, following the setting in [20]. All tasks are fine-tuned using sparse rewards, which provide feedback only in the form of success or failure signals.
Environments: Google Research Football. To evaluate NCDPO’s capability in large discrete action spaces, we test on three Google Research Football scenarios requiring multi-agent coordination: 3 vs 1 with Keeper, Counterattack Hard, and Corner. Here we adopt a centralized control strategy, where actions for all agents are generated simultaneously using a single Diffusion Policy. The base Diffusion Policies are pre-trained to output one-hot vectors corresponding to ground-truth actions. To construct the pre-training dataset, we aggregate trajectories from multiple MLP-based policies with varying success rates.
6.2 Evaluation on Continuous Robot Control Tasks
We first evaluate NCDPO on continuous control tasks across two benchmarks: OpenAI Gym locomotion and Robomimic. In these environments, we compare NCDPO with DPPO [20]; DRWR and DAWR [20], based on reward-weighted regression [17] and advantage-weighted regression respectively [16]; DIPO [31], which employs action gradients as the score function for denoising steps; and Q-learning-based methods such as IDQL [8] and DQL [28].
From the experimental results shown in Fig. 3 and Fig. 4, we observe that NCDPO consistently achieves the strongest performance and exhibits robustness across all tasks. While DPPO, which is the best among the baseline methods, performs comparably to NCDPO in the Robomimic benchmark  [^1] Note that we tested performance using the latest version of DPPO (v0.8), which is about 2.5x sample efficient in Transport task as reported in the original paper., it lags behind in the OpenAI Gym locomotion environments. Other baselines generally underperform relative to DPPO. Notably, IDQL demonstrates strong performance on the first three Robomimic tasks but fails in the final one. In contrast, DQL suffers from instability across all scenarios. We additionally conducted experiments on Square using vision-based inputs, with results provided in the appendix. The results demonstrate consistent improvements in Diffusion Policy performance when fine-tuned with NCDPO.
Figure 3: Performance comparison on OpenAI Gym locomotion tasks. Results are averaged over three seeds. NCDPO (ours) achieves the strongest performance.
Figure 4: Performance comparison on Robomimic tasks. Results are averaged over three seeds. NCDPO (ours) achieves the strongest performance.
Scenario NCDPO (Ours) DPPO AWR IDQL DQL RWR DIPO
Hopper-Medium 126.4 (1.8) 98.4 (2.0) 44.8 (1.2) 113.8 (0.2) 122.7 (1.2) 100.9 (3.6) 94.4 (4.9)
Hopper-Medium-Replay 128.2 (2.8) 114.6 (3.3) 82.1 (7.0) 117.9 (1.1) 121.1 (3.6) 104.2 (3.1) 108.6 (1.0)
Hopper-Medium-Expert 135.2 (2.0) 102.4 (5.2) 40.7 (0.5) 131.9 (0.4) 120.9 (28.1) 113.6 (4.7) 127.5 (1.3)
Walker2d-Medium 118.4 (3.8) 101.2 (1.6) 93.5 (8.3) 110.7 (0.5) 94.9 (36.9) 90.2 (3.3) 99.8 (2.0)
Walker2d-Medium-Replay 126.6 (4.5) 105.1 (4.3) 75.8 (2.3) 121.9 (2.6) 107.2 (37.8) 69.2 (6.0) 88.6 (11.0)
Walker2d-Medium-Expert 141.0 (2.3) 137.5 (2.0) 124.2 (5.4) 135.5 (2.7) 67.3 (47.6) 106.8 (6.8) 138.8 (1.4)
HalfCheetah-Medium 122.0 (11.0) 82.3 (0.7) 65.5 (2.9) 79.3 (0.8) 77.1 (5.4) 47.9 (5.4) 73.9 (0.6)
HalfCheetah-Medium-Expert 139.7 (6.8) 80.6 (1.6) 64.4 (2.3) 77.8 (1.0) 72.1 (8.0) 38.4 (2.9) 71.4 (0.2)
HalfCheetah-Medium-Replay 121.0 (2.0) 72.3 (0.4) 60.5 (0.7) 74.3 (0.9) 73.0 (3.1) 30.7 (1.8) 58.1 (0.9)
Lift 100.0 (0.0) 99.7 (0.2) 93.3 (1.7) 99.2 (0.1) 99.8 (0.3) 97.5 (0.5) 97.3 (0.8)
Can 99.3 (1.2) 99.0 (1.0) 33.8 (3.2) 94.5 (3.1) 0.3 (0.6) 90.7 (0.8) 52.8 (5.1)
Square 87.3 (4.5) 87.0 (2.3) 40.3 (8.5) 80.0 (5.0) 0.0 (0.0) 74.8 (2.5) 25.3 (4.5)
Transport 96.7 (2.31) 91.3 (2.9) 11.2 (3.5) 0.5 (0.8) 0.0 (0.0) 0.0 (0.0) 0.2 (0.3)
Table 1: Mean and standard deviation of performance over continuous robot control scenarios. Each result is evaluated on three different seeds. NCDPO (ours) exhibits the strongest performance. Performance on OpenAI Gym locomotion tasks are normalized according to scores of MLP policies trained from scratch using PPO with 1M samples reported in Tianshou [29] . Original scores are listed in Table 3.
6.3 Evaluation on Discrete Multi-agent Coordination Tasks
[…]
As Fig. 5 demonstrates, NCDPO outperforms the MLP baseline across all three evaluated scenarios. This outcome not only highlights the superiority of Diffusion Policy in handling complex and diverse demonstration data over simple MLP policy, but also confirms the effectiveness of the DP+NCDPO during the fine-tuning phase.
Figure 5: Performance comparison in Google Research Football. Results are averaged over at least three seeds. NCDPO (ours) exhibits strong performance and stability.
Scenario NCDPO (Ours) MAPPO
3 vs 1 with Keeper 87.4(2.7) 75.1(12.3)
Counterattack Hard 87.0(2.5) 80.0(2.0)
Corner 78.3(4.5) 74.9(3.0)
Table 2: Average evaluation success rate and standard deviation (over three seeds) on Google Research Football scenarios. The base Diffusion Policy and MLP policy are pre-trained on the same dataset. MLP policy is trained using Cross-Entropy loss.
6.4 NCDPO is Robust to the Number of Denoising Steps
We further conduct an ablation study to investigate the impact of varying the number of denoising steps in the diffusion model. The experimental results shown in Fig. 7 indicate that NCDPO demonstrates strong robustness to the choice of denoising steps.
Figure 6: Ablation Study on Denoising Steps in OpenAI Gym locomotion tasks.
Figure 7: Ablation Study on Denoising Steps in Robomimic tasks.
We hypothesize that the robustness of NCDPO arises from the way gradients are propagated through time during the diffusion process. This gradient flow leads to more accurate gradient estimates.
7 Conclusion and Limitations
We present NCDPO, a novel approach for fine-tuning Diffusion Policies through Proximal Policy Optimization that exhibits strong performance across continuous and discrete control domains. Our key innovation lies in reformulating the diffusion denoising process as a noise-conditioned stochastic policy that enables effective gradient backpropagation through diffusion timesteps. Through extensive experiments across locomotion, manipulation, and multi-agent cooperation scenarios, we demonstrate that NCDPO achieves superior sample efficiency and final performance compared to existing diffusion RL approaches. NCDPO’s ability to handle both continuous and discrete action spaces suggests its potential as a general-purpose policy optimization framework.
Our study focuses on the algorithmic development and evaluation of NCDPO in simulated settings. Consequently, we have not yet explored sim-to-real transfer on physical robots. These choices reflect our emphasis on fine-tuning methodology. Extending NCDPO to real-world deployment remains to be implemented in future work.
Appendix A Self-Imitation Regularizer
When directly fine-tuning Diffusion Policies using policy gradient methods, we observe a structure collapse issue—namely, the Diffusion Policy fails to maintain consistency between the forward and reverse processes. To preserve the structural integrity of the diffusion model, we introduce self-imitation regularization. Specifically, we perform behavior cloning on the trajectories generated in the previous episode. Empirically, we find that this regularization significantly reduces the behavior cloning loss. In contrast, without it, this behavior cloning loss will keep increasing, indicating structural degradation in the Diffusion Policy.
Appendix B Additional experimental results
B.1 Original Scores on OpenAI Gym locomotion tasks
Scenario NCDPO (Ours) DPPO AWR IDQL DQL RWR DIPO PPO
Hopper-Medium 3297.3 (47.8) 2566.6 (51.1) 1168.9 (30.5) 2970.2 (5.2) 3200.7 (30.1) 2633.6 (94.0) 2463.1 (127.6) 2609.3
Hopper-Medium-Replay 3345.64 (71.88) 2988.97 (86.90) 2142.30 (183.82) 3076.91 (29.82) 3159.72 (94.85) 2718.27 (79.86) 2834.07 (25.00) 2609.3
Hopper-Medium-Expert 3528.74 (51.89) 2672.64 (135.15) 1062.47 (13.15) 3440.50 (10.79) 3153.66 (733.36) 2964.11 (121.35) 3326.99 (32.73) 2609.3
Walker2d-Medium 4248.8 (137.6) 3632.1 (55.9) 3353.9 (296.9) 3972.8 (17.3) 3405.2 (1322.7) 3238.3 (116.8) 3581.6 (70.3) 3588.5
Walker2d-Medium-Replay 4544.59 (162.98) 3770.52 (154.50) 2719.04 (83.84) 4373.89 (91.66) 3846.26 (1357.97) 2483.04 (215.70) 3180.80 (396.27) 3588.5
Walker2d-Medium-Expert 5060.92 (82.87) 4935.57 (73.28) 4458.68 (195.26) 4863.91 (95.34) 2416.68 (1708.34) 3831.65 (243.76) 4979.96 (50.22) 3588.5
HalfCheetah 7058.8 (635.1) 4758.3 (41.8) 3788.7 (166.5) 4584.4 (45.3) 4459.1 (309.5) 2773.1 (310.0) 4272.4 (33.2) 5783.9
HalfCheetah-Expert 8079.30 (392.21) 4663.23 (92.62) 3723.21 (131.37) 4499.39 (57.61) 4171.97 (465.25) 2218.39 (168.36) 4126.96 (12.14) 5783.9
HalfCheetah-Replay 7000.14 (113.88) 4181.57 (24.59) 3501.08 (40.34) 4295.81 (53.02) 4223.07 (177.28) 1775.98 (101.28) 3362.48 (49.47) 5783.9
Table 3: Mean and standard deviation of original scores across continuous robot control scenarios.
B.2 Experiments with Vision Inputs
We performed evaluation on Square task in robomimic with vision as input. Results demonstrate the effectiveness of NCDPO.
Figure 8: Experimental results for vision inputs.
B.3 Ablation Study
We observe that setting the action chunk size to one significantly improves performance in Gym environments. We hypothesize that this is due to the nature of these tasks, where agents must respond promptly to rapid and continuous changes in the environment. Smaller chunk sizes allow the policy to adapt its actions more frequently, which is crucial for achieving fine-grained control. The corresponding results are presented in Figure 9.
Figure 9: Ablation study on action chunk size in OpenAI Gym locomotion tasks.
additionally, in discrete action environments, modifying the noise scheduler to increase Gaussian noise during the denoising steps improves exploration without degrading overall performance, as shown in Figure 10. In discrete settings, the absolute values of the logits are less important than their relative magnitudes, which allows increased noise to encourage exploration while preserving policy effectiveness. To achieve this, we adjust the noise scheduler using parameters $\eta$ and $\beta_{\text{base}}$, increasing the noise level via the transformation:
$\beta^{\prime}_{k}=\beta_{base}\left(\frac{\beta_{k}}{\beta_{base}}\right)^{\eta}$
where $\beta_{k}$ corresponds to the original noise schedule defined in Equation 3. In our implementation, we set $\beta_{base}=0.7$.
Figure 10: Ablation study on different values of $\eta$ in Google Research Football.
We further find that increasing the initial noise scale $\sigma_{a}$ in the acting layer enhances exploration. An ablation study conducted on Robomimic supports this finding:
Figure 11: Ablation study of different choices of initial $\log\sigma_{a}$ in Robomimic tasks.
B.4 Further Experiments in OpenAI Gym locomotion tasks.
We evaluated different training methods using datasets of varying quality for pretraining the base policy. The "medium-replay" dataset consists of replay buffer samples collected before early stopping, while the "medium-expert" dataset contains equal proportions of expert demonstrations and suboptimal rollouts [6].
Regardless of dataset quality, NCDPO consistently outperforms all baselines, as shown in Figures 12 and 13.
Figure 12: Pretraining with expert datasets.
Figure 13: Pretraining with replay datasets.
B.5 Google Research Football Data Curation
For each scenario, we collected 200K environment steps per model. The win rates of the agents used for dataset generation are summarized in Table 4.
Scenario Win Rates
3 vs 1 with Keeper 0.93, 0.90, 0.70, 0.55
Corner 0.76, 0.75, 0.50, 0.50, 0.41
Counterattack Hard 0.90, 0.78, 0.70, 0.61, 0.56, 0.56
Table 4: Win rates of trained agents used for dataset collection in Google Research Football. Each model contributes 200,000 steps.
Appendix C Implementation Details and Hyperparameters
For NCDPO, we apply Adam optimizer for actor and AdamW optimizer for critic. For all other baselines, AdamW optimizer is adopted.
For fair comparison, we adopt the same network architecture as DPPO [20] and directly utilize their implementation for model structure. Our overall training framework is built upon a modified codebase of MAPPO [32].
Task $\gamma$ $\lambda$ Action Chunk Actor LR Critic LR Actor MLP Size Critic MLP Size Actor MLP Layers $\eta$ Initial Noise Log Std $1/T$ Denoising Step Clone Epochs Clone LR Episode Length Mini-batch Number Environment Max Steps Parallel Environments
Hopper 0.995 0.985 4 3e-5 1e-3 1024 256 7 0 -2 - 5 8 1e-3 256 1 1000 32
Walker2d 0.995 0.985 4 3e-5 1e-3 1024 256 7 0 -2 - 5 8 1e-3 500 1 1000 32
HalfCheetah 0.99 0.985 4 3e-5 1e-3 1024 256 7 0 -2 - 5 3 1e-4 500 1 1000 32
lift 0.999 0.99 4 3e-5 1e-3 1024 256 7 0 -2.3 - 5 2 1e-3 300 4 300 200
can 0.999 0.99 4 3e-5 1e-3 1024 256 7 0 -2.3 - 5 60 3e-4 300 4 300 200
square 0.999 0.99 4 3e-5 1e-3 1024 256 7 0 -2.3 - 5 200 2e-4 400 4 400 200
square-vision 0.999 0.99 4 3e-5 5e-4 1024 256 7 0 -2.3 - 5 200 2e-4 400 4 400 200
transport 0.999 0.99 8 3e-5 1e-3 1024 256 7 0 -2.3 - 5 321 8e-4 800 4 800 200
3 vs 1 with Keeper 0.99 0.95 1 3e-5 1e-3 1024 256 7 0.03 - 20 5 10 1e-3 200 1 - 50
Corner 0.99 0.95 1 3e-5 1e-3 1024 256 7 0.03 - 20 5 10 1e-3 500 1 - 50
Counterattack Hard 0.99 0.95 1 3e-5 1e-3 1024 256 7 0.03 - 20 5 10 1e-3 500 1 - 50
Table 5: Hyperparameter settings for different tasks of NCDPO.
Task $\gamma$ $\lambda$ Action Chunk Actor LR Critic LR Actor MLP Size Critic MLP Size Actor MLP Layers Denoising Step Episode Length Mini-batch Size Environment Max Steps Parallel Environments
Hopper 0.99 0.95 4 1e-4 1e-3 512 256 3 20 2000 50000 1000 40
Walker2d 0.99 0.95 4 1e-4 1e-3 512 256 3 20 2000 50000 1000 40
HalfCheetah 0.99 0.95 4 1e-4 1e-3 512 256 3 20 2000 50000 1000 40
lift 0.999 0.95 4 1e-4 5e-4 512 256 3 20 1200 7500 300 50
can 0.999 0.95 4 1e-4 5e-4 512 256 3 20 1200 7500 300 50
square 0.999 0.95 4 1e-4 5e-4 512 256 3 20 1600 10000 400 50
transport 0.999 0.95 8 1e-4 5e-4 512 256 3 20 3200 10000 800 50
Table 6: Hyperparameter settings for Baselines in robot control. Experiment is executed using DPPO [20] implementation and hyperparameters. Batch size for all baselines other than DPPO is 1000. For further details, please refer to DPPO paper [20].
Task $\gamma$ $\lambda$ Action Chunk Actor LR Critic LR Actor MLP Size Critic MLP Size Actor MLP Layers $\eta$ Initial Noise Log Std $1/T$ Denoising Step Clone Epochs Clone LR
3 vs 1 with Keeper 0.99 0.95 1 5e-4 5e-4 256 256 - - - 20 5 8 1e-3
Corner 0.99 0.95 1 5e-4 5e-4 256 256 - - - 20 5 8 1e-3
Counterattack Hard 0.99 0.95 1 5e-4 5e-4 256 256 - - - 20 5 8 1e-3
Table 7: Hyperparameter of MLP on football. Experiment is run on MAPPO codebase and MLP architecture remains same as MAPPO, and does not use residual connection, thus rendering parameter MLP layers unusable.
Task $\gamma$ $\lambda$ Action Chunk Actor LR Critic LR Actor MLP Size Critic MLP Size Actor MLP Layers $\eta$ Initial Log Std $1/T$ Denoising Step
Walker2d-NCDPO 0.995 0.985 1 1e-4 1e-3 256 256 3 0 -0.8 - 5
HalfCheetah-NCDPO 0.99 0.985 1 1e-4 1e-3 256 256 3 0 -0.8 - 5
Walker2d-MLP+PPO 0.995 0.985 1 1e-4 1e-3 256 256 3 - -0.8 - -
HalfCheetah-MLP+PPO 0.99 0.985 1 1e-4 1e-3 256 256 3 - -0.8 - -
Walker2d-DPPO 0.99 0.985 1 1e-4 1e-3 512 256 3 - - - 10
HalfCheetah-DPPO 0.99 0.985 1 1e-4 1e-3 512 256 3 - - - 10
Table 8: Hyperparameters for training from scratch. In this experiment, MLP+PPO has exatcly the same architecture with MLP in diffusion’s denoising process. Numbers of mini-batches and parallel environments are the same as Table 5.
Task $\gamma$ $\lambda$ Action Chunk Actor LR Critic LR Actor MLP Size Critic MLP Size Actor MLP Layers $\eta$ Initial Log Std $1/T$ Denoising Step Clone Epochs Clone LR
Hopper 0.995 0.985 4 3e-5 1e-3 1024 256 7 0 -2 - 5/10/20 8 1e-3
Walker2d 0.995 0.985 4 3e-5 1e-3 1024 256 7 0 -2 - 5/10/20 8 1e-3
HalfCheetah 0.99 0.985 4 3e-5 1e-3 1024 256 7 0 -2 - 5/10/20 8 1e-3
square 0.995 0.985 4 3e-5 5e-4 1024 256 7 0 -2.3 - 5/20 8 5e-4
transport 0.99 0.985 8 3e-5 5e-4 1024 256 7 0 -2.3 - 5/20 8 5e-4
Table 9: Hyperparameter of Ablation on Denoising Steps.
Task $\gamma$ $\lambda$ Action Chunk Actor LR Critic LR Actor MLP Size Critic MLP Size Actor MLP Layers $\eta$ Initial Log Std $1/T$ Denoising Step Clone Epochs Clone LR
lfit 0.999 0.99 4 3e-5 1e-3 1024 256 7 0 -2.3/-2.7 - 5 2 1e-3
can 0.999 0.99 4 3e-5 1e-3 1024 256 7 0 -2.3/-2.7 - 5 60 3e-4
square 0.999 0.99 4 3e-5 1e-3 1024 256 7 0 -2.3/-2.7 - 5 200 2e-4/5e-4
transport 0.999 0.99 8 3e-5 1e-3 1024 256 7 0 -2.3/-2.7 - 5 321 8e-4/5e-4
Table 10: Hyperparameter of Ablation on Initial Noise.
Task $\gamma$ $\lambda$ Action Chunk Actor LR Critic LR Actor MLP Size Critic MLP Size Actor MLP Layers $\eta$ Initial Log Std $1/T$ Denoising Step Clone Epochs Clone LR
3 vs 1 with Keeper 0.99 0.95 1 3e-5 1e-3 1024 256 7 0.03/0.1/0.3/1 - 20 5 8 1e-3
Corner 0.99 0.95 1 3e-5 1e-3 1024 256 7 0.03/0.1/0.3/1 - 20 5 8 1e-3
Counterattack Hard 0.99 0.95 1 3e-5 1e-3 1024 256 7 0.03/0.1/0.3/1 - 20 5 8 1e-3
Table 11: Hyperparameter of Ablation on $\eta$. $\eta=1$ indicates using original scheduler.
Appendix D Computational Resources
Each run could be done in 6 hours with 1 AMD Ryzen 3990X 64-Core Processor and 1 NVIDIA 3090 GPU.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our main contribution, a novel framework for Diffusion Policy fine-tuning is properly described in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations are mentioned in the conclusion.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not have theory included.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have clearly described the algorithm in the main text and listed all necessary details in the appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have included the details and will release code and datasets soon.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have listed details in appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We averaged results over at least 3 different seeds and reports mean and standard deviation. Meanwhile, the performance advantage is significant enough.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We have provided compute resources in appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, we followed the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our paper focus on designing a novel training framework, and no societal impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our paper has no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Datasets are cited in main text, code is cited in Appendix C
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We have included our code and dataset in supplementary materials.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLM is only used for paper language refinement and code auto-completion.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.